OpenStudy (anonymous):

PreCal- So my formula is 17512=27(2t+1.2)^2-113(2t+1.2)+86 For some reason no matter what I do I'm getting different answers and wrong answers. PLEASE help.

OpenStudy (amistre64):

have you tried doing it to a little lynyrd skynyrd?

OpenStudy (anonymous):

{{t == -12.299157316325548}, {t == 13.19174990891814}}

OpenStudy (anonymous):

I think t=13.19175, -12.29916? Not positive though

OpenStudy (amistre64):

substitue a variable to clean things up... u = (2t+1.2)^2 17512= 27u^2 -113u +86 make it equal to zero 0 = 27u^2 -113u +86 - 17512

OpenStudy (amistre64):

0 = 27u^2 -113u +17426

OpenStudy (amistre64):

anything we can factor out to begin with?

OpenStudy (amistre64):

if anything, itll setup like this.... but you might need to just use the quadratic formula on it or work thru completeing the square (u- /27)(u- /27)

OpenStudy (amistre64):

yeah, i miss typed a (+)... :)

OpenStudy (anonymous):

The number of bacteria in a refrigerated food product is given by , where is the temperature of the food. When the food is removed from the refrigerator, the temperature is given by , where is the time in hours. Hence the formula I got which is correct. NOW i have to ... Find the time when the bacteria count reaches 17512 which is only 1 answer in the form of hours.

OpenStudy (amistre64):

u = 27.9691 and u = -23.0431 which one do you wanna use?

OpenStudy (amistre64):

ack..... use googles calculator :)

OpenStudy (anonymous):

LOL ok so I didn't think of the fact one was negative. Obviously I don't have negative hours.

OpenStudy (amistre64):

then well use u = 35,090.5556 2t+1.2 = 35,090.5556 t = [35,090.5556 -1.2] / 2 if im lucky lol

OpenStudy (amistre64):

17 544.6778 maybe??

OpenStudy (anonymous):

13.19174990891814 this is the answer. I have no idea how sambol and mmbuckaroos got it though. I distributed first, then foiled the first one and then factored. 0= 2916t^2+3273.2t-16240.64 and couldn't solve it from there. So I tried substituting x for (2t+1.2) So.. 27x^2-113x-17426=0 and still nothing...